A ball is thrown vertically upwards from a platform 8 feet above the ground. After 2 seconds, the

ball is 2 feet below the platform. Find the initial velocity of the ball.


1
Expert's answer
2021-04-26T02:04:00-0400
"y=h+vt-0.5gt^2\\2=8+2v-0.5(32.174)(2)^2\\v=29.2\frac{ft}{s}"